OpenStudy (anonymous):

-3x^2-8x+1=-10x using quadratic formula can someone explain it step by step im so confused

mathslover (mathslover):

Hello @sciencewiz123 : First of all transpose -10 x from Right Hand Side to Left Hand Side : \[\large{-3x^2 - 8 x + 1 + 10x = 0}\] \[\large{-3x^2 - 8x + 10x + 1 =0}\] \[\large{-3x^2 + 2x + 1 =0}\] ^ This is what we have as in the form of quad. equation now. Solve this by suing quadratic formula given below : \[\large{\textbf{root of the equation } = \frac{-b \pm \sqrt{b^2-4ac}}{2a}}\] Where b is here 2 a is here (-3) and c = 1

mathslover (mathslover):

Do not forget to post here the answer what you get after solving it, also if you get any problem understanding this^ , post here.